Evaluate sin-(sin (57)).
Answer #1

sin (sin (ST) = sin(sin ( 17 - 17/18 ) ) - sin (coz I ) = sin(-0-984807753) = - sint(0.784807753) IY [:siñ(-x) = -sin(x)]